# Mitigating Graph Covariate Shift via Score-based Out-of-distribution Augmentation

### v1

@@ A.4.2 Implementation settings
$s_{\theta}$ Number of graph transformer layers $8$ $8$ $9$ $8$
Number of attention heads $8$ $8$ $8$ $8$
Hidden dimension of $\bm{X}$ $256$ $256$ $256$ $256$
Hidden dimension of $\bm{A}$ $64$ $64$ $64$ $64$
SDE for $\bm{X}$ Type VP VP VP VP
@@
Weight decay $1\times 10^{-12}$ $1\times 10^{-12}$ $1\times 10^{-12}$ $1\times 10^{-12}$
@@
Graph Out-of-Distribution Classification: Following prior work (Gui et al., 2022; Li et al., 2024), we employ GIN-Virtual (Xu et al., 2018; Gilmer et al., 2017) as the GNN backbone for the GOOD-CMNIST, GOOD-HIV, and GOOD-SST2 datasets. For the GOOD-Motif dataset, we adopt GIN (Xu et al., 2019). To ensure a fair comparison across all methods, we utilize the same GNN backbone architecture for all models.
@@
We perform a grid search for the hyperparameter $\alpha\in\{0.5,1.0\}$ across all datasets and find that $\alpha=0.5$ consistently yields satisfactory results throughout the experiments. For $\lambda$, the grid search is tailored to each dataset. Specifically, we explore $\lambda\in\{0.01,0.02,0.03,0.04,0.05\}$ for the GOOD-Motif-base and GOOD-HIV-scaffold datasets. For GOOD-CMNIST-color, we tune $\lambda\in\{0.05,0.1\}$. In the case of GOOD-SST2-length, where $\lambda=0.01$ corresponds to an increase of one node in the graph size relative to the training distribution, we expand the grid search to $\lambda\in\{0.01,0.02,...,0.14\}$. Similarly, for GOOD-Motif-size, where $\lambda=0.01$ reflects an increase of one node, we use a search space of $\lambda\in\{0.01,0.02,0.03,0.04,0.05\}$. For GOOD-HIV-size, where $\lambda=0.01$ corresponds to a decrease of ten nodes in graph size from the training distribution, we also use $\lambda\in\{0.01,0.02,0.03,0.04,0.05\}$. Since this hyperparameter tuning is performed during the sampling phase rather than the training phase, it is not computationally intensive.
A.5 Baseline settings
@@
The implementation details for GNN backbones and hyperparameter tuning are consistent with those outlined in Appendix A.4.2. For methods including ERM, IRM (Arjovsky et al., 2019), GroupDRO (Sagawa et al., 2019), VREx (Krueger et al., 2021), DANN (Ganin et al., 2016), Deep Coral (Sun & Saenko, 2016), DIR (Wu et al., 2022), DropNode (Feng et al., 2020), DropEdge (Rong et al., 2019), MaskFeature (Thakoor et al., 2021), FLAG (Kong et al., 2022), M-Mixup (Wang et al., 2021), and G-Mixup (Han et al., 2022), we report results from the study in (Li et al., 2024), which uses the same GNN backbones and hyperparameter tuning as specified in Appendix A.4.2. For GSAT (Miao et al., 2022), CIGA (Chen et al., 2022), and AIA (Sui et al., 2024), we use their publicly available source code, adopting default settings and hyperparameters as detailed in their papers.
A.6 Additional Experimental Results
We also utilize the GOOD-SST2-length dataset to validate that OODA can generate an augmented distribution, $\tilde{P}_{\mathrm{tr}}(G,Y)$, in a controlled manner while preserving stable patterns. In the context of GOOD-SST2-length, the parameter $\lambda$ regulates the OOD size of the augmented graphs, with $\lambda=0.1$ corresponding to an increase of one unit in the graph size from the training distribution. We systematically vary $\lambda$ within the range $[0,1)$ in increments of 0.1 to generate ten augmented datasets, each containing the same number of graphs as the training set. The results for GOOD-SST2-length under length covariate shift are illustrated in Figure 5. As shown in Figure 5 (Left), the MMD RBF between the original GOOD-SST2-length graph distribution, $P_{\mathrm{tr}}(G,Y)$, and the augmented graph distribution, $\tilde{P}_{\mathrm{tr}}(G,Y)$, increases with increasing values of $\lambda$. Specifically, when $\lambda=0.0$, the MMD RBF is $0.032\pm 0.005$. As $\lambda$ rises to $0.9$, the MMD RBF escalates to $25$ times its value at $\lambda=0.0$. These results indicate that OODA can effectively generate OOD graph samples in a controlled manner by flexibly adjusting $\lambda$.
Moreover, as depicted in Figure 5 (Right), the expected probability $\mathbb{E}_{\tilde{G_{0}}\sim\tilde{P}}\left[p\left(\mathbf{y}_{G}\mid\tilde{G_{0}}\right)\right]$ of the augmented GOOD-SST2-length graph distribution consistently exceeds $0.94$ as $\lambda$ increases. This result demonstrates the capability of OODA to generate OOD graphs that retain stable patterns.
Figure 5: (Left): Distance between the original GOOD-SST2-length graph distribution $P_{\mathrm{tr}}(G,Y)$ and the augmented graph distribution $\tilde{P}_{\mathrm{tr}}(G,Y)$. (Right): Expected probabilities that the augmented GOOD-SST2-length graph distributions retain stable patterns that determine the target class.

### v2

@@ D.2. Implementation settings
$s_{\theta}$ Number of graph transformer layers 88 88 99 88
Number of attention heads 88 88 88 88
Hidden dimension of $\boldsymbol{X}$ $256$ $256$ $256$ $256$
Hidden dimension of $\boldsymbol{A}$ $64$ $64$ $64$ $64$
SDE for $\boldsymbol{X}$ Type VP VP VP VP
@@
Weight decay $\times 10^{-12} 1×1$$\times 10^{-12}$ $\times 10^{-12} 1×1$$\times 10^{-12}$
@@
Graph Out-of-Distribution Classification: Following prior work (gui2022good; ligraph), we employ GIN-Virtual (xu2018powerful; gilmer2017neural) as the GNN backbone for the GOOD-CMNIST, GOOD-HIV, and GOOD-SST2 datasets. For the GOOD-Motif dataset, we adopt GIN (xu2018how). To ensure a fair comparison across all methods, we utilize the same GNN backbone architecture for all models.
@@
We perform a grid search for the hyperparameter $\alpha\in\{0.1,0.5,1.0\}$ across all datasets. For $\lambda$, the grid search is tailored to each dataset. Specifically, we explore $\lambda\in\{0.01,0.02,0.03,0.04,0.05\}$ for the GOOD-Motif-base dataset and $\lambda\in\{0.1,0.2,0.3\}$ for GOOD-HIV-scaffold dataset. For GOOD-CMNIST-color, we tune $\lambda\in\{0.05,0.1\}$. In the case of GOOD-SST2-length, where $\lambda=0.01$ corresponds to an increase of one node in the graph size relative to the training distribution, we expand the grid search to $\lambda\in\{0.01,0.02,...,0.14\}$. Similarly, for GOOD-Motif-size, where $\lambda=0.01$ reflects an increase of one node, we use a search space of $\lambda\in\{0.01,0.02,0.03,0.04,0.05\}$. For GOOD-HIV-size, where $\lambda=0.01$ corresponds to a decrease of ten nodes in graph size from the training distribution, we also use $\lambda\in\{0.01,0.02,0.03,0.04,0.05\}$. Since this hyperparameter tuning is performed during the sampling phase rather than the training phase, it is not computationally intensive.
Appendix E Baseline settings
@@
The implementation details for GNN backbones and hyperparameter tuning are consistent with those outlined in Appendix D.2. For methods including ERM, IRM (arjovsky2019invariant), GroupDRO (sagawa2019distributionally), VREx (krueger2021out), DANN (ganin2016domain), Deep Coral (sun2016deep), DIR (wu2022discovering), DropNode (feng2020graph), DropEdge (rong2019dropedge), MaskFeature (thakoor2021large), FLAG (kong2022robust), M-Mixup (wang2021mixup), and G-Mixup (han2022g), we report results from the study in (ligraph), which uses the same GNN backbones and hyperparameter tuning as specified in Appendix D.2. For GSAT (miao2022interpretable), CIGA (chen2022learning), and AIA (sui2024unleashing), we use their publicly available source code, adopting default settings and hyperparameters as detailed in their papers.
Appendix F Sensitivity of hyperparameter $\lambda$
We determine the value of $\lambda$ by evaluating its effectiveness on OOD validation sets across various datasets. The sensitivity of our method to different $\lambda$ values is illustrated in Tables 6, 7 and 8.
Table 6. Performance on the GOOD-Motif-base dataset across varying $\lambda$ values.
$\lambda$ GOOD-Motif-base
0.01 91.80
0.02 92.97
0.03 92.83
0.04 93.03
0.05 92.77
Table 7. Performance on the GOOD-HIV-scaffold dataset across varying $\lambda$ values.
$\lambda$ GOOD-HIV-scaffold
0.1 78.94
0.2 77.73
0.3 77.42
Table 8. Performance on the GOOD-CMNIST-color dataset across varying $\lambda$ values.
$\lambda$ GOOD-CMNIST-color
0.05 68.66
0.1 67.91
Appendix G Statistical Significance
Statistical significance is evaluated via a paired nonparametric bootstrap test (davison1997bootstrap) at the $5\%$ level. The observed gains are significant across settings: on GOOD-Motif (base), CODA improves over DropNode by $0.70\%$ with $p=0.0305$; on GOOD-Motif (size), over AIA by $4.96\%$ with $p=$ 0.0015 ; on GOOD-CMNIST (color), over AIA by $18.23\%$ with $p=4.9998\times 10^{-5}$; on GOOD-HIV (scaffold), over DropNode by $1.49\%$ with $p=0.0147$; on GOOD-HIV (size), over DANN by $1.28\%$ with $p=0.0062$; and on GOOD-SST2 (length), over MaskFeature by $0.69\%$ with $p=0.0207$.
Appendix H Time and memory complexity
Our pipeline consists of two stages: data augmentation, and trainng of the GNN classifier on augmented graphs. The second stage of classification follows general GNN training setup, without introducing additional complexity: the time/memory complexity per layer of using Graph Isomorphism Networks (GIN) as backbone is $\Theta(n+e)$, where $n$ is the number of nodes and $e$ is the number of edges. For the data augmentation stage, we introduce graph transformer whose memory and time complexity per layer is $\Theta(n^{2})$. This arises from the computation of attention scores and predictions across each edge.